OpenStudy (asdfghj8):

Solve for v. 5v^2 +5=-26v

Vocaloid (vocaloid):

re-arrange to the form ax^2 + bx + c = 0 5v^2 + 26v + 5 = 0 then factor

Vocaloid (vocaloid):

(5v +1)(v+5) = 0 solve for v

OpenStudy (asdfghj8):

so it would be v=-5 and v=-1/5 ??

Vocaloid (vocaloid):

yes, good job!

Vocaloid (vocaloid):

idk where the other person got their answers from

OpenStudy (asdfghj8):

Thank you so much for explaining and helping me understand how you got it. I really wanted to know the process lol